2 Direct Labor Variances Bellingham Company produces a product that requires 4 standard hours per unit...

2

Direct Labor Variances

Bellingham Company produces a product that requires 4 standard hours per unit at a standard hourly rate of $16.00 per hour. If 4,800 units required 19,800 hours at an hourly rate of $15.36 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) total direct labor cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $
b. Direct labor time variance $
c. Total direct labor cost variance $
Answer #1

Labor rate variance = (SR-AR) * AH

= (16 - 15.36) * 19,800

= - 12,672 (Minus sign here) f

Labor efficiency variance = (SH-AH) * SR

= (4800*4 - 19,800) * 16

= 9600

Total labor variance

= - 12,672 + 9600

= - 3072

Comment if you face any issues